# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ 3.1 An integral representation of the free energy
$A^{\prime}(\beta)$ $=$ $\frac{1}{Z_{\beta,\pi}(\mathbf{x})}\int_{\mathcal{H}}\hat{L}(h,\mathbf{x})e^{-\beta\hat{L}(h,\mathbf{x})}d\pi(h)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x}}\big[\hat{L}(h,\mathbf{x})\big],$
$A^{\prime\prime}(\beta)$ $=$ $-\Big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})^{2}\big]-\big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]\big)^{2}\Big)\leq 0\text{.}$
The equality in the lemma then follows from the first two identities above and the fundamental theorem of calculus, and the inequality follows from the last identity, which shows that $\mathbb{E}_{g\sim G_{\beta_{k-1},\pi}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$ is non-increasing in $\beta$. ∎
In statistical physics there is a formal analogy, where the function $h\mapsto\hat{L}(h,\mathbf{x})$ is the ($\mathbf{x}$-dependent) energy of the system in the state $h$, and $\beta$ is the inverse temperature. The Gibbs posterior then becomes the "canonical ensemble" (Gibbs, 1902), describing the probability of states in equilibrium with a heat bath at temperature $\beta^{-1}$. The function $\beta\mapsto A(\beta)$ plays an important role: $\beta^{-1}A(\beta)$ is the Helmholz free energy, $A^{\prime}(\beta)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]$ is the thermal average of the energy, $\beta A^{\prime}(\beta)+A(\beta)$ is the entropy and $A^{\prime\prime}(\beta)$ is proportional to the heat capacity at temperature $\beta^{-1}$ (see e.g. Huang, 2008).
@@
For $h\in\mathcal{H}$, $\mathbf{x\in}\mathcal{X}^{n}$ and an increasing sequence $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ of positive numbers, we denote
$\Gamma(h,\mathbf{x},\bm{\beta})=-\beta_{K}\hat{L}(h,\mathbf{x})+\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$ (3)
So Lemma 3.1 states that, when $\beta_{0}=0$, then,
$\beta_{K}\hat{L}\left(h,\mathbf{x}\right)-\ln Z_{\beta_{K}}\left(\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right).$ (4)
Note that $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends only on the training data $\mathbf{x}$, the sequence $\bm{\beta}$ and the hypothesis $h$.
3.2 Bounds
The function $F$ in the following is a placeholder for a random variable related to the generalization gap, which we would like to bound with high probability.
@@ 3.2 Bounds
Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function, $\beta>0$ and $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ as above with $\beta_{0}=0$ and $\beta_{K}=\beta$. Then,
(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$

### v2

@@ 3 Bounds for the Gibbs Algorithm
$A^{\prime}(\beta)$ $=$ $\frac{1}{Z_{\beta,\pi}(\mathbf{x})}\int_{\mathcal{H}}\hat{L}(h,\mathbf{x})e^{-\beta\hat{L}(h,\mathbf{x})}d\pi(h)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big],$
$A^{\prime\prime}(\beta)$ $=$ $-\text{Var}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]\leq 0,$
where Var denotes variance. (3) then follows from the first two of these identities and the fundamental theorem of calculus, and the last assertion in the lemma follows from the last identity, since zero variance at any $\beta$ implies that $\hat{L}\left(h,\mathbf{x}\right)$ is almost surely constant with respect to $\pi$. Since the logarithm of the density of the Gibbs posterior is $\beta\hat{L}\left(.,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)$ we get (4). Then (5) follows from taking the expectation of (4) in $G_{\beta}\left(\mathbf{x}\right)$. ∎
Figure 3: The mean training error of the Gibbs posterior is plotted against the inverse temperature. The relative entropy $KL\left(G_{\beta}\left(\mathbf{x}\right),\pi\right)$ is equal to the shaded area AA.
Figure 3 provides a simple geometrical interpretation for conclusion (5). For conclusion (4), if the shaded area were equal to $\ln(dG_{\beta}\left(\mathbf{x}\right)/d\pi)\left(h\right)$, its lower boundary would fluctuate with a variance equal to the negative slope at $\beta$, depending on the draw of $h$ (compare the proof of Lemma 3.1).
Equations (4) and (5) of Lemma 3.1 can now be substituted into Theorem 2.1 to give our bounds for the Gibbs algorithm.
4 Bounds for Langevin Monte Carlo
The Gibbs posterior is an idealization, from which it is impossible to sample directly. Nevertheless, a number of works describe Markov processes, here summarized as Langevin Monte Carlo (LMC), capable of approximating a probability measure $\nu$ on $\mathbb{R}^{d}$ of the form $\nu\propto\exp\left(-V\right)$, or some nearby limiting distribution.
The classical prototype is Continuous Langevin Dynamics (CLD), a Markov process in continuous time describing thermalization in statistical physics and originating in the study of Brownian motion (Langevin and others, 1908). To turn the continuous process into an iterative algorithm several discretized versions have been proposed. There is the Metropolis Adjusted Langevin Algorithm (MALA) (Roberts and Tweedie, 1996), which is the Euler-discretization of CLD with an additional Metropolis-style accept-reject step to ensure that the invariant distribution is indeed the desired Gibbs distribution. Hamiltonian Monte Carlo Markov Chain (HMCMC) (Duane et al., 1987; Betancourt, 2017) is a refinement of MALA. The Unadjusted Langevin Algorithm (ULA) (Parisi, 1981) is the discretization of CLD without the accept-reject step and converges to a slightly different distribution. Stochastic Gradient Langevin Dynamics (SGLD) (Gelfand and Mitter, 1991; Welling and Teh, 2011) is an accelerated version of ULA replacing the true gradient of $V$ by an unbiased estimate realized with minibatches. All these processes are of theoretical interest as models for Stochastic Gradient Descent (SGD), but at least SGLD is also used as a learning algorithm in practice.
In Appendix C.1, we give more recent references containing convergence guarantees and discuss CLD and ULA in some detail. In the next section we show how the PAC-Bayesian bound and our integral representation can be applied to time-homogeneous Markov processes.
Throughout this section, we assume $\mathcal{H}=\mathbb{R}^{d}$ and an isotropic Gaussian prior $\pi$ of width $\sigma$. We condition on the training data $\mathbf{x}$, reference to which we often omit.
4.1 Bounds for Markov processes
In this section, we both apply and improve the method in the recent work of Harel et al. (2025) on Markov processes. We take $\left\{h_{t}\left(\mathbf{x}\right)\right\}_{t\in I}$ to be a time-homogeneous Markov-process in real or discrete time, $I=\left[0,\infty\right)$ or $I=\mathbb{N}_{0}$, with values in $\mathcal{H}$. This is our model of a training process, such as CLD or the practically implementable MALA, ULA or SGLD. The distribution of $h_{t}\left(\mathbf{x}\right)$ will be denoted $\nu_{t}\left(\mathbf{x}\right)$. We will use the following result, sometimes referred to as the second law of thermodynamics. A proof is given in Section C.3.
Lemma 4.1.
(Second law of thermodynamics) If $\nu$ is a stationary distribution of $\left\{h_{t}\right\}_{t\in I}$ and $s<t$ then $KL\left(\nu_{t},\nu\right)\leq KL\left(\nu_{s},\nu\right)$ and $R_{\infty}\left(\nu_{t},\nu\right)\leq R_{\infty}\left(\nu_{s},\nu\right)$, with equality in either case if and only if $\nu_{s}=\nu$.
Now for any $t\in I$ and any $\beta>0$
$KL\left(\nu_{t},\pi\right)$ $=$ $\mathbb{E}_{h\sim\nu_{t}}\left[\ln\frac{d\nu_{t}}{dG_{\beta}}\right]+\mathbb{E}_{h\sim\nu_{t}}\left[\ln\frac{dG_{\beta}}{d\pi}\right]$ (6)
$=$ $KL\left(\nu_{t},G_{\beta}\right)-\beta~\mathbb{E}_{h\sim\nu_{t}}\left[\hat{L}\left(h\right)\right]-\ln Z_{\beta}.$
There is an analogous identity for the single draw, omitting the expectations in $\nu_{t}$ and replacing $KL$ by $R_{\infty}$. Now assume that $G_{\beta}$ is a stationary distribution of the process, and that $\nu_{0}=\pi$, so the process is started from the prior. Then, Harel et al. (2025) use Lemma 4.1 above, to get $KL\left(\nu_{t},G_{\beta}\right)\leq KL\left(\pi,G_{\beta}\right)=\beta\ \mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]+\ln Z_{\beta}$, and they substitute this bound in the above identity. Since the partition functions cancel each other, and since $\beta\mathbb{E}_{h\sim\nu_{t}}\left[\hat{L}\left(h\right)\right]\geq 0$, they end up with $KL\left(\nu_{t},\pi\right)\leq\beta\ \mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$, to be substituted in the PAC-Bayesian bound. There is a similar bound for the single draw in terms of the relative Rényi entropy $R_{\infty}\left(\nu_{\beta,t},G_{\beta}\right)$. The resulting generalization bounds of Harel et al. (2025) remain valid along the entire training trajectory. They are, however, largely independent of distribution and data, and vacuous for $\beta>n$, since typically $\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$and $\left\|\hat{L}\right\|_{\infty}$are on the order of unity or larger (here, we omitted several substantial refinements in Harel et al. (2025)).
The use of the second law is elegant, but in this form it forgoes the potential benefit of a process converging to $G_{\beta}$ (as CLD, MALA, or HMCMC) or to some nearby distribution (as with ULA or SGLD). The following proposition takes advantage of convergence as well as of close approximation.
Proposition 4.2.
Assume $\nu_{0}=\pi$. (i) Let $\lambda_{t}=R_{\infty}\left(\nu_{t},G_{\beta}\right)/R_{\infty}\left(\pi,G_{\beta}\right)$. Then
$R_{\infty}\left(\nu_{t},\pi\right)\leq\lambda_{t}\beta~\left\|\hat{L}\right\|_{\infty}+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
(ii) If instead $\lambda_{t}=KL\left(\nu_{t},G_{\beta}\right)/KL\left(\pi,G_{\beta}\right)$, then
$KL\left(\nu_{t},\pi\right)\leq\lambda_{t}\beta~\mathbb{E}_{g\sim\pi}\left[\hat{L}\left(g\right)\right]+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
Proof.
We only prove (i), the proof of (ii) being analogous.
$R_{\infty}\left(\nu_{t},\pi\right)$ $=$ $\sup_{h\in\mathcal{H}}\left(\ln\frac{d\nu_{t}}{dG_{\beta}}\left(h\right)+\ln\frac{dG_{\beta}}{d\pi}\left(h\right)\right)$
$\leq$ $\lambda_{t}R_{\infty}\left(\pi,G_{\beta}\right)+\sup_{h\in\mathcal{H}}\left(-\beta\hat{L}\left(h\right)-\ln Z_{\beta}\right)$
$\leq$ $\lambda_{t}\sup_{h\in\mathcal{H}}\left(\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right)-\ln Z_{\beta}$
$=$ $\lambda_{t}\beta~\left\|\hat{L}\right\|_{\infty}+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma,$
where we used Lemma 3.1 in the last step. ∎
Remarks: 1. Without any additional assumptions these bounds remain valid for all $\beta$ and along the entire training trajectory. For all $t$, where $\lambda_{t}\leq 1$, the bounding expression is a convex combination. By the last assertion of Lemma 3.1 the integral is generically smaller than $\beta~\left\|\hat{L}\right\|_{\infty}$ or $\beta~\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$ respectively. So whenever $\lambda_{t}<1$ our bounds are also smaller than those in Section 3 of Harel et al. (2025). For ULA and SGLD we expect that $\lambda_{t}$ converges to some value $\lambda_{\infty}\ll 1$.
2. If $G_{\beta}$ is indeed stationary, then by the second law (Lemma 4.1) $\lambda_{t}$ is non-increasing in $t$ and $\lambda_{t}<1$ for $t>0$ in all non-trivial cases. The bounds then move in convex interpolation towards the integral.
3. If the process converges to $G_{\beta}$ in relative entropy, meaning that $KL\left(\nu_{t},G_{\beta}\right)\rightarrow 0$, then $\lambda_{t}\rightarrow 0$ as $t\rightarrow\infty$. In this case the bounds asymptotically approach those of the Gibbs posterior in Lemma 3.1, with the modification that the negative terms are omitted. For large $t$ they exhibit the same sensitivity of generalization to mean training errors of the Gibbs posterior at higher temperatures. This is the case for CLD and all algorithms containing a Metropolis-style accept-reject mechanism, such as MALA or HMCMC.
To illustrate this, assume that $\mathcal{H}=\mathbb{R}^{d}$ with Gaussian prior $\pi$ of width $\sigma$. When adapting the stochastic differential equation for CLD to temperature and prior, it becomes
$dh_{\beta,t}=-\left(\nabla\hat{L}\left(h_{\beta,t}\right)+\frac{h_{\beta,t}}{\beta\sigma^{2}}\right)dt+\sqrt{\frac{2}{\beta}}dB_{t},$
where $B_{t}$ is standard centered Brownian motion in $\mathbb{R}^{d}$. Let $\nu_{\beta,t}$ be the distribution of $h_{\beta,t}$ at time $t$. In Lemma C.3 in Appendix C.2 we show, that, if $G_{\beta}$ satisfies a logarithmic Sobolev inequality with constant $\alpha$ (see Section C.1), then $KL\left(\nu_{\beta,t},G_{\beta}\right)\leq e^{-2\alpha t/\beta}KL\left(\nu_{\beta,0},G_{\beta}\right)$. So if CLD starts from the prior, we can use Proposition 4.2 (ii) with $\lambda_{t}=e^{-2\alpha t/\beta}$ and obtain
$KL\left(\nu_{\beta,t},\pi\right)\leq e^{-2\alpha t/\beta}~\beta~\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]+\left(1-e^{-2\alpha t/\beta}\right)\int_{0}^{\beta}\mathbb{E}_{G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
4.2 Stability of the Bounds
If $\nu:\mathcal{X}^{n}\to[0,\infty)$ is the stochastic algorithm for which we want the bound, then the results of the previous section show, that with sufficient approximation of $\nu\left(\mathbf{x}\right)$ by $G_{\beta}\left(\mathbf{x}\right)$ in relative entropy we can largely recover the bounds for the Gibbs posterior. But these bounds, though data-dependent, are still inaccessible because of the continuous nature of the temperature integral and the impossibility to sample directly from the Gibbs posterior. We now study the following question of stability: given some algorithm to approximate $G_{\beta}$ for any $\beta$ to arbitrary precision in relative entropy, can we also approximate our bounds to arbitrary precision?
To this end we discretize the temperature scale of the integral and approximate each $G_{\beta_{k}}$ by some distribution $\nu_{k}$. The error incurred on the corresponding expectations of $\hat{L}$, can then be controlled under either boundedness or Lipschitz conditions on the loss $\ell$.
Definition 4.3.
@@ 4.2 Stability of the Bounds
For $\mathbf{x\in}\mathcal{X}^{n}$, an increasing sequence $\beta_{0}^{K}=\left(0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta\right)$ of positive numbers, and a corresponding vector of data-dependent distributions $\nu_{0}^{K-1}(\mathbf{x})=\left(\nu_{0}\left(\mathbf{x}\right),\nu_{1}\left(\mathbf{x}\right),\cdots,\nu_{K-1}\left(\mathbf{x}\right)\right)\in\mathcal{P}\left(\mathcal{H}\right)^{K}$ we denote
$\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})=\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim\nu_{k-1}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$
The next lemma bounds the estimation error relative to the temperature integral in terms of the relative entropies.
Lemma 4.4.
With $\beta_{0}^{K}$ and $\nu_{0}^{K-1}$ as in Definition 4.3 denote
$\Delta:=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]d\gamma-\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$
(i) If $\mathbb{E}_{h\sim G_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim\nu_{k}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$ for all $k$ and $\mathbf{x}$, then $\Delta\leq 0$.
(ii) If $\ell\left(h,\mathbf{x}\right)$ is bounded in $h$ for all $\mathbf{x}$, $\left\|\ell\left(h,\mathbf{x}\right)\right\|\leq m$ then
$\Delta\leq m\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\sqrt{KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)/2}.$
(iii) If instead $\ell\left(h,\mathbf{x}\right)$ is $m$-Lipschitz in $h$ for all $\mathbf{x}$, $\ell\left(h,\mathbf{x}\right)-\ell\left(g,\mathbf{x}\right)\leq m\left\|h-g\right\|$ and $G_{\beta_{k}}\left(\mathbf{x}\right)$ satisfies an LSI with constant $\alpha$ for all $k$ and $\mathbf{x}$, then
$\Delta\leq\frac{2m}{\alpha}\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1}){KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)}.$
By the last conclusion of Lemma 3.1 part (i) is immediate. Proofs of (ii) and (iii) are given in Appendix C.4. The assumption in case (i) is not implausible, if we start LMC from a non-informative prior, and in our experiments we always observed decreasing losses along the LMC path.
The next theorem gives our final bound in terms of arbitrary distributions and their relative entropies to Gibbs distributions.
@@
Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function and $\beta_{0}^{K}$ and $\nu_{0}^{K-1}$ as in Definition 4.3. Let $\nu(\mathbf{x})$ be any data-dependent distribution on $\mathcal{H}$. Let $\Delta$ be bounded as in Lemma 4.4, depending on which of the conditions there is fulfilled by $\ell$. Then
(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu(\mathbf{x})$